# Notions of enriched purity

### v2

@@ 9. The canonical language
Remark 9.3.
If the factorization system is proper then $\mathcal{M}$ contains all the regular monomorphisms (see the dual of [9, 2.1.4])). Thus $\parallel\phi\parallel_{M}\cong M_{f,g}$.
@@
where $\widehat{\mathcal{K}}(f,g,K)$ is the pullback of $\mathcal{K}(f,K)$ along $\mathcal{K}(g,K)$ and $h$ is the morphism induced into the product. Since $\mathcal{E}$ is closed under composition, the interpretation of a formula of the form $\psi(x)\equiv\exists y\ (s_{f}(x)=s_{g}(y))$ is given by the $(\mathcal{E},\mathcal{M})$-factorization of the map
@@
where $\parallel\psi\parallel_{K}$ and $\parallel\psi\parallel_{L}$ are by definition the $(\mathcal{E},\mathcal{M})$-factorizations of $\mathcal{K}(g,K)$ and $\mathcal{K}(g,L)$ respectively, and the middle vertical arrow is induced by the factorization. Then by definition we have $\parallel\psi\parallel_{L}\cong\mathcal{P}(g,L)$ as $\mathcal{M}$-subobjects of $\mathcal{K}(A,L)$. Now, $f$ is $(\lambda,\mathcal{E})$-pure with respect to $g$ if and only if the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$ is given by $\mathcal{P}(g,f)$, if and only if $\parallel\psi\parallel_{K}\cong\mathcal{P}(g,f)$ as $\mathcal{M}$-subobjects of $\mathcal{K}(A,K)$, if and only if $\parallel\psi\parallel_{K}$ is the pullback of $\parallel\psi\parallel_{L}$ along $\mathcal{K}(A,f)$, if and only if $f$ is elementary with respect to $\psi$.
@@
involving an identity morphism, meaning that $\parallel\phi\parallel_{X}\cong\parallel\phi^{\prime}\parallel_{X}$. Given $\phi$ as above, with $g\colon C\to A$ and $h\colon C\to B$, let $D$ be the pushout of $g$ along $h$ in $\mathcal{K}$, together with the induced morphism $h^{\prime}\colon A\to D$. Then $D$ is still $\lambda$-presentable and $\phi$ is equivalent in $\mathcal{K}$ to the formula $\phi^{\prime}$ with the $h^{\prime}$ just defined. This is thanks to Remark 9.4 and the fact that $\widehat{\mathcal{K}}(g,h,X)\cong\mathcal{K}(D,X)$ for any $X\in\mathcal{K}$. (2) Let $(\mathcal{E},\mathcal{M})$ be proper, and consider now a general pp-formula
@@
meaning that $\parallel\phi_{1}\wedge\cdots\wedge\phi_{n}\parallel_{X}\cong\parallel\phi\parallel_{X}$ as $\mathcal{M}$-subobjects of $\mathcal{K}(A,X)\times\mathcal{K}(B,X)$, for any $X\in\mathcal{K}$. Note that, by Remark 9.3, $\parallel\phi_{i}\parallel_{X}$ is the pullback of $\mathcal{K}(g_{i},X)$ along $\mathcal{K}(h_{i},X)$; thus $\parallel\phi_{1}\wedge\cdots\wedge\phi_{n}\parallel_{X}$, being obtained by pulling back all of the $\parallel\phi_{i}\parallel_{X}$, coincides with the pullback of $\mathcal{K}(g,X)$ along $\mathcal{K}(h,X)$, which is just $\parallel\phi\parallel_{X}$. Now the result follows from point (1). ∎

### v3

@@ Appendix A The canonical language
Remark A.3.
If the factorization system is proper then $\mathcal{M}$ contains all the regular monomorphisms (see the dual of [10, 2.1.4])). Thus $\parallel\phi\parallel_{M}\cong M_{f,g}$.
@@
where $\widehat{\mathcal{K}}(f,g,K)$ is the pullback of $\mathcal{K}(f,K)$ along $\mathcal{K}(g,K)$ and $h$ is the morphism induced into the product.
Since $\mathcal{E}$ is closed under composition, the interpretation of a formula of the form $\psi(x)\equiv\exists y\ (s_{f}(x)=s_{g}(y))$ is given by the $(\mathcal{E},\mathcal{M})$-factorization of the map
@@
where $\parallel\psi\parallel_{K}$ and $\parallel\psi\parallel_{L}$ are by definition the $(\mathcal{E},\mathcal{M})$-factorizations of $\mathcal{K}(g,K)$ and $\mathcal{K}(g,L)$ respectively, and the middle vertical arrow is induced by the factorization. Then by definition we have $\parallel\psi\parallel_{L}\cong\mathcal{P}(g,L)$ as $\mathcal{M}$-subobjects of $\mathcal{K}(A,L)$.
Now, $f$ is $(\lambda,\mathcal{E})$-pure with respect to $g$ if and only if the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$ is given by $\mathcal{P}(g,f)$, if and only if $\parallel\psi\parallel_{K}\cong\mathcal{P}(g,f)$ as $\mathcal{M}$-subobjects of $\mathcal{K}(A,K)$, if and only if $\parallel\psi\parallel_{K}$ is the pullback of $\parallel\psi\parallel_{L}$ along $\mathcal{K}(A,f)$, if and only if $f$ is elementary with respect to $\psi$.
@@
involving an identity morphism, meaning that $\parallel\phi\parallel_{X}\cong\parallel\phi^{\prime}\parallel_{X}$. Given $\phi$ as above, with $g\colon C\to A$ and $h\colon C\to B$, let $D$ be the pushout of $g$ along $h$ in $\mathcal{K}$, together with the induced morphism $h^{\prime}\colon A\to D$. Then $D$ is still $\lambda$-presentable and $\phi$ is equivalent in $\mathcal{K}$ to the formula $\phi^{\prime}$ with the $h^{\prime}$ just defined. This is thanks to Remark A.4 and the fact that $\widehat{\mathcal{K}}(g,h,X)\cong\mathcal{K}(D,X)$ for any $X\in\mathcal{K}$.
(2) Let $(\mathcal{E},\mathcal{M})$ be proper, and consider now a general pp-formula
@@
meaning that $\parallel\phi_{1}\wedge\cdots\wedge\phi_{n}\parallel_{X}\cong\parallel\phi\parallel_{X}$ as $\mathcal{M}$-subobjects of $\mathcal{K}(A,X)\times\mathcal{K}(B,X)$, for any $X\in\mathcal{K}$. Note that, by Remark A.3, $\parallel\phi_{i}\parallel_{X}$ is the pullback of $\mathcal{K}(g_{i},X)$ along $\mathcal{K}(h_{i},X)$; thus $\parallel\phi_{1}\wedge\cdots\wedge\phi_{n}\parallel_{X}$, being obtained by pulling back all of the $\parallel\phi_{i}\parallel_{X}$, coincides with the pullback of $\mathcal{K}(g,X)$ along $\mathcal{K}(h,X)$, which is just $\parallel\phi\parallel_{X}$. Now the result follows from point (1). ∎